OpenStudy (anonymous):

write an equation in standard form for the line that is perpendicular to the graph of 3y-4x=1 and has a y intercept of 4

OpenStudy (anonymous):

please help i need this fast!

Nnesha (nnesha):

ok solve for y first

Nnesha (nnesha):

do you know how to solve ??????????/

OpenStudy (anonymous):

no

Nnesha (nnesha):

ok no problem so add 4x both side and let me know what you got

OpenStudy (anonymous):

3y= 1+4x

Nnesha (nnesha):

divide by 3

OpenStudy (anonymous):

y= 1/3+ 4/3

Nnesha (nnesha):

you need equation that is perpendicular to other equation and you already have y-intercept which is 4

OpenStudy (anonymous):

so y= 1/3+ 1/4x?

Nnesha (nnesha):

so when you need to change to perpendicular just flip the slope

OpenStudy (anonymous):

i am so confused

OpenStudy (anonymous):

so y= 1/3+ 2x is the answer?

Nnesha (nnesha):

how you get 2x and remember y-intercept of new equation is 4

OpenStudy (anonymous):

do you know what the answer is?

Nnesha (nnesha):

 you should do like this